### v2

# ITEACNet: Inverted Teacher-studEnt seArch Conversation Network

Abstract
Due to the widespread applications of conversations in human-computer interaction, Conversation Emotion Recognition (CER) has attracted increasing attention from researchers. In real-world scenarios, the emotional states of both participants in a conversation tend to maintain a relatively stable pattern within the local context, and often encountering issues with incomplete data patterns. Focusing on these two key challenges, we propose a novel framework for incomplete multimodal learning in CER, called ”Inverted Teacher-studEnt seArch Conversation Network (ITEACNet).” ITEACNet comprises two novel components: the ”Emotion Context Changing Encoder (ECCE)” and the ”Inverted Teacher-Student framework (ITS).” ECCE considers context changes from both local and global perspectives, while the ITS allows a simple teacher model to learn complete data processing methods, enabling a complex student model to follow the performance of the teacher model using incomplete data. Furthermore, we employ a Neural Architecture Search algorithm to enhance the capabilities of student model , achieving superior model performance. Finally, to align with real-world scenarios, we introduce a novel evaluation method, testing the model’s performance under different missing rate conditions without altering the model weights. We conduct experiments on three benchmark CER datasets, and the results demonstrate that our ITEACNet outperforms existing methods in incomplete multimodal CER.
*Equal contribution
Introduction
Conversation Emotion Recognition (CER) aims to analyze the emotional states of speakers in conversational contexts. It has found extensive applications in areas such as social media analysis and human-computer interaction (Lian et al. 2023). Recent studies have achieved promising results in CER. However, there are still some limitations that need to be addressed.
First, unlike sentence-level emotion understanding tasks, conversations feature longer sequences and important context information. Long-sequence conversations often entail complex context changes, the model need to carefully consider these changes to better understand conversation data. Second, real-world scenarios introduce various challenges that impact the collection of multimodal data. Factors such as background noise in speech, obstacles in images, and recognition errors in text may result in the omission of modal information, introducing uncertainty to the completeness of the data. Consequently, the model must exhibit robustness in addressing incomplete data.
In this paper, we propose a novel framework for incomplete learning in CER, called ”Inverted Teacher-studEnt seArch Conversation Network (ITEACNet)”. Fig 1 illustrates the overall structure of our method. To better understand the context changes in conversations, we design a novel encoding method that enables the model to simultaneously consider them from both local and global perspectives, called Emotion Context Change Encoding (ECCE). To enhance the model’s ability to handle missing data, we propose a novel training framework, called Inverted Teacher-Student (ITE), which allows the model to utilize computationally intensive techniques, such as Neural Architecture Search (NAS) (Liu, Simonyan, and Yang 2019), to more fully utilize undamaged data to impute missing data.
To verify the effectiveness of our method, we conduct experiments on three benchmark CER datasets. However, due to the uncertain nature of data incompleteness in real-world scenarios, existing evaluation methods, which involve training model weights separately for testing under various missing rate scenarios (Sun et al. 2023; Lian et al. 2023; Peng et al. 2020), fail to authentically reflect the model’s robustness. Consequently, we propose a novel evaluation method to simulate more realistic missing scenarios. Specifically, we fix the model weights and assess its performance across different missing rates, providing a more accurate measure of the model’s robustness.
Through quantitative and qualitative analysis, we demonstrate that our ITEACNet outperforms currently advanced approaches. The main contributions of this paper are summarized as follows:
  • We design a novel encoding method, called Emotion Context Change Encoding, that enables the model to simultaneously consider the context changes from both local and global perspectives.
  • We propose a novel Inverted Teacher-Student training method as the foundational framework for incomplete data, and propose the utilizing of NAS to enhance the imputation of the student model.
  • Unlike existing works that train individual models for specific missing rates, we propose a novel evaluation approach, namely assessing the average performance of a unified model across all missing rates. This evaluation method is closer to the real-world scenario and fills the gap in current research.
  • Experimental results on three benchmark datasets demonstrate that our method, ITEACNet, is superior to existing state-of-the-art approaches in CER.
Figure 1: The overall structure of Inverted Teacher-studEnt seArch Conversation Network (ITEACNet) with the trimodal setting. This model focuses on the Conversation Emotion Recognition (CER) task based on incomplete data.
Figure 2: Encoder Layer calculation.
Related Work
Learning from incomplete multimodal data is a fundamental research area in machine learning. One simple approach involves data imputation, followed by utilizing existing classification methods on the imputed data. Adding zero vectors or using the mean values of the same modality within a class as missing data are the most straightforward methods (Parthasarathy and Sundaram 2020; Zhang et al. 2020, 2022). However, the performance of such approaches inevitably reaches an upper limit, as they cannot leverage deeper multimodal correlations to improve downstream task performance. Consequently, some studies start leveraging the correlations between different modalities to optimize this process, manifesting as a low-rank data matrix resulting from high correlation (Yang et al. 2017; Liang et al. 2019). Many works, therefore, project the data into a common space using low-rankness (Fan et al. 2017; Liang et al. 2019). Nevertheless, for large data, these methods often incur high computational costs.
With the advancement of deep learning, deep models have achieved breakthroughs in the generation domain, and numerous deep learning-based imputation methods have demonstrated superior capabilities. For instance, Duan et al. (Duan et al. 2014) utilize autoencoders to estimate missing data, achieving commendable results in incomplete multimodal learning. Tran et al.(Tran et al. 2017) further propose the Cascaded Residual Autoencoder (CRA) to enhance the imputation capability of autoencoders. Additionally, Zhao et al. (Zhao, Li, and Jin 2021) combine CRA with cycle-consistency loss for cross-modal imputation, outperforming existing methods. Furthermore, Zuo et al. (Zuo et al. 2023) and Sun et al. (Sun et al. 2023) enhance the imputation ability of autoencoders for missing data from both unimodal and multimodal perspectives, or both high- and low-level features, achieving outstanding performance. Meanwhile, Lian et al. (Lian et al. 2023) delve deeper into considering temporal and speaker information in conversations, achieving state-of-the-art performance in CER.
Certainly, besides imputation, there exist alternative strategies for handling incomplete data. For instance, some approaches partition incomplete data into multiple subsets with complete modalities to model them independently (Yuan et al. 2012; Li et al. 2018). Simultaneously, some works enforce low-dimensional representations for different modalities to ensure high correlation within the same sample (Hotelling 1992; Andrew et al. 2013). However, this work mainly focuses on interpolation methods, so it will not be introduced much.
Method
Input Format
In this task, a conversation $C={\{(u_{i},y_{i})\}}^{L}_{i=1}$ consists of $L$ utterances, and the $i^{th}$ utterance $u_{i}$ with a label $y_{i}$. For each utterance, we extract features $x_{m}\in\mathbb{R}^{1\times d_{m}}$, $m\in\{t,a,v\}$, from text, audio and visual modalities, where $d$ represents the feature dimension, and concatenate them in the order of speaking to obtain $X_{m}\in\mathbb{R}^{L\times d_{m}}$, to represent the entire conversation.
Inverted Teacher-Student
Given that current imputation strategies for missing data mainly focus on the vector distance between complete and incomplete data, dealing with incomplete data introduces an additional imputation task alongside the recognition task for the model. This necessitates more complex computations compared to tasks involving only recognition. Therefore, we design a Inverted Teacher-Student training framework, where two different models learn two types of data. The teacher model requires sufficient recognition capability with complete data, while the student model employs a more complex structure to learn how to process incomplete data.
Teacher-Student training framework is designed to distill knowledge from a complex teacher model into a simpler student model, Inverted Teacher-Student serves as a framework allowing a complex student to emulate the performance of a simple teacher. Our framework is illustrated in Figure 1.
We define complete multimodal features as $\{X_{t},X_{a},X_{v}\}$ and incomplete multimodal data as $\{X_{t}^{s},X_{a}^{s},X_{v}^{s}\}$. We first compute their Emotion Context Change Encoding (ECCE):
$\{E_{t},E_{a},E_{v}\}=\mathrm{ECCEEncoder}(\{X_{t},X_{a},X_{v}\})$ (1)
$\{E_{t}^{s},E_{a}^{s},E_{v}^{s}\}=\mathrm{ECCEEncoder_{s}}(\{X_{t}^{s},X_{a}^{s},X_{v}^{s}\})$ (2)
where $E_{m}\in\mathbb{R}^{L\times L}$. Subsequently, they pass through multiple Encoder Layers in their respective models, generating hidden features under the guidance of ECCE:
$\{H_{i}\}_{i=1}^{L_{tea}}=\mathrm{Teacher}(\{X_{t},X_{a},X_{v}\},\{E_{t},E_{a},E_{v}\})$ (3)
$\{H_{j}^{s}\}_{j=1}^{L_{stu}}=\mathrm{Student}(\{X_{t}^{s},X_{a}^{s},X_{v}^{s}\},\{E_{t}^{s},E_{a}^{s},E_{v}^{s}\})$ (4)
$L_{tea}$ and $L_{stu}$ denote the numbers of Encoder Layers for the Teacher Model and Student Model, respectively. It’s worth noting that, for ease of understanding, we label the hidden feature closest to the classification layer as $H_{1}$. To enhance the computing capability of the Student Model, we allocate more Encoder Layers to it.
In the end, the final layer of hidden features $H_{1}$ and $H_{1}^{s}$ is employed by the classifier to predict emotion labels. Simultaneously, the masked portions of the last three layers of hidden features from the Student Model, $\{H_{1}^{s},H_{2}^{s},H_{3}^{s}\}$, are computed for imputation loss, aiming to closely approximate the vector of the Teacher Model, $\{H_{1},H_{2},H_{3}\}$.
Emotion Context Changing Encoder
Figure 3: In conversations, the speakers’ states tend to maintain a relatively stable pattern within the local context. As the context change, this pattern evolves.
As depicted in Figure 3, the emotional state of speakers in a conversation tends to maintain a stable pattern locally, which evolves with context changes. To enhance model’s perception of these dynamics, we design a novel encoding method, Emotion Context Correlation Encoding. The encoder consists of two specific steps: encoding local emotion context and encoding global context changing.
Encoding local emotion context
Given a context window size $w$, for the $i^{th}$ utterance, the calculation process for its local context feature $z^{i}_{m}$ can be described as follows:
$z^{i}_{m}=F_{\text{local}}^{m}(X_{m}\left[i-\lceil\frac{w}{2}\rceil:i+\lfloor\frac{w}{2}\rfloor\right])$ (5)
$F_{\text{local}}^{m}$ represents a replaceable function used for processing features within the window. After calculating for each utterance, we obtain the local context feature $Z_{m}\in\mathbb{R}^{L\times d_{z}}$.
Encoding global context changing
The local context feature considers only the emotion context within a limited window. However, to account for the dynamics of the context changing over time, further calculations are needed for the $Z_{m}$. For the context changing feature $e^{i,j}_{m}$ from the $i^{th}$ to $j^{th}$ utterances, the calculation process can be described as follows:
$e^{i,j}_{m}=F_{\text{global}}^{m}(Z_{m}\left[i:j\right])$ (6)
$F_{\text{global}}^{m}$ represents a replaceable function used for processing context features. After calculating for each utterance pair, we obtain the context changing feature $E_{m}\in\mathbb{R}^{L\times L\times d_{e}}$.
Algorithm 1 Pseudo code of ECCE Encoder.
Input: $X_{m}\in\mathbb{R}^{L\times d_{m}}$, $w$, $dist_{m}$
Output: $E_{m}\in\mathbb{R}^{N\times N\times d_{e}}$
$F_{\text{local}}^{m}$ = Conv1d($d_{m}$, $d_{z}$, $w$, $pool=w//2$, $stride=1$)
$F_{\text{global}}^{m}$ = Mean()
$FC$ = Linear($d_{z},d_{e}$)
[…]
$E_{m}$ = Zeros($N,N,d_{z}$)
  for $i$ in $range(0,N)$ do
[…]
        $E_{m}[i][j]$ = $F_{\text{global}}^{m}$($Z_{m}[i:j]$)
        $E_{m}[j][i]$ = $E_{m}[i][j]$
     end for
  end for
$E_{m}$ = $FC$($E_{m}$)
Return: $E_{m}$
Algorithm 1 presents the pseudocode for the Emotion Context Changing Encoder. In our experiments, we use a simple 1D convolution operation as $F_{\text{local}}$ and a straightforward average pooling operation as $F_{\text{global}}$.
ITEACNet
To enhance the model’s perception of emotions, we leverage ECCE to guide the modeling of single-modal sequence information. Additionally, to improve the model’s imputation capability for incomplete data, we propose the use of NAS to enhance student. In this section, we will provide a detailed explanation of these processes. As illustrated in Figure 2, the Encoder Layer and ITEAC Encoder subgraphs depict the model we designed, while the Token Mixer subgraph illustrates the guidance process of ECCE.
Encoder Layer
As shown in Figure 2, multimodal features undergo encoding in their respective ITEAC Encoder at each Encoder Layer. Subsequently, the Cross Attention module merges various modal information into text features. The fused text features will serve as the latent vector $H$ in Figure 1. Subsequently, the three modal data will continue to be fed into the next Encoder Layer. The calculation process of Cross Attention can be described as follows:
$output=\text{softmax}(\frac{qk^{T}}{\sqrt{C}})v$ (7)
$C$ represents the dimension of $q$. In both rounds of Cross Attention modules, text features consistently serve as $q$, while audio and visual features respectively function as $k$ and $v$ for the two iterations.
ITEAC Encoder
In the design of the ITEAC Encoder, we retain the overall framework of Metaformer (Yu et al. 2022b) and conduct a search for the Token Mixer (Yang et al. 2022). Unlike previous workes, to maximize the model’s encoding capacity, we utilize the Router module (Yu et al. 2022a) to extend the architecture search process to input-dependent. Different inputs can choose different architectures as needed, rather than opting for a single architecture for all inputs. Given the candidate module set $\mathbb{O}$, where $\mathbb{|O|}$ represents the number of modules, the search process through the Router can be described as follows:
$X^{m}_{\alpha}=\text{softmax}(\text{Router}(X_{m}))$ (8)
$output=\sum_{o\in\mathbb{O}}{\sum_{i=1}^{L}{X^{m}_{\alpha}\left[:,i\right]\times o(X_{m}\left[i\right])}}$ (9)
where $X^{m}_{\alpha}\in\mathbb{R}^{L\times|\mathbb{O}|}$ and $output\in\mathbb{R}^{L\times d_{m}}$.
While there have been numerous excellent designs for Token Mixers recently, our focus is on the introduction of NAS methods rather than algorithm design. Therefore, we select 6 simple and lightweight operations as candidates. Please refer to the appendix for a detailed description of them.
In our experiments, when the Router is a linear layer and the Token Mixer comprises the selected six candidates, we denote the ITEAC Encoder as NAS. Conversely, when the Router is an Identity operation and the Token Mixer includes only Self-Attention, we denote it as Transformer.
Guidance of ECCE
The role of the Token Mixer is to interact with information from different utterances in conversation. A category of mixers, represented by attention operation, typically generates a Mix Map, which is normalized and used as weights for fusing utterance feature. The emotion context change information we designed serves to adjust the Mix Map, thereby guiding the interaction of sequential information (Ying et al. 2021). The specific calculation process can be described as follows:
$MixMap_{e}=MixMap+E_{m}$ (10)
$output=X_{m}\otimes\mathrm{softmax}(MixMap_{e})$ (11)
Joint Optimization
To simultaneously train the model for both the target task and handling incomplete data, we jointly optimize the emotion loss, $L_{emo}$, and the imputation loss, $L_{rec}$.
Consistent with prior work, we employ cross-entropy loss as the emotion loss for labeled category data. The calculation formula is shown as follows:
$L_{emo}=-\frac{1}{2L}\sum_{i=1}^{L}{(y_{i}log(\hat{y}_{i}^{t})+y_{i}log(\hat{y}_{i}^{s}))}$ (12)
$y_{i}\in\mathbb{R}^{y}$ is the true label. For labeled dimensional value data, we use the mean square error as the emotion loss. The predicted values for this data are transformed into category labels based on a domain range conversion before evaluating classification performance. The loss calculation formula is shown as follows:
$L_{emo}=\frac{1}{2L}(||y_{i}-\hat{y}_{i}^{t}||^{2}+||y_{i}-\hat{y}_{i}^{s}||^{2})$ (13)
To enable the student model to better handle incomplete data, we simultaneously calculate the reconstruction loss for its last three layers of hidden features. We assign a progressively weighted adjustment to the model’s attention to each layer’s reconstruction loss:
$L_{rec}=0.5||H_{1}^{s}-H_{1}||^{2}+0.1||H_{2}^{s}-H_{2}||^{2}+0.05||H_{3}^{s}-H_{3}||^{2}$ (14)
The hidden layer vectors of the teacher model will be treated as constants during the calculation; therefore, the reconstruction loss will not backpropagate any gradients to the teacher model.
Finally, we merge these two loss functions into a unified objective function. This combined loss is employed to optimize all trainable parameters in an end-to-end manner.
$Loss=L_{emo}+L_{rec}$ (15)
Experiments
Table 1: Ablation results for Emotion Context Change Encoding.
Dataset Local Global WAF
IEMOCAP (four-class) 77.83
77.69
78.15
IEMOCAP (six-class) 60.28
59.89
60.66
CMU-MOSI 85.46
85.16
85.74
CMU-MOSEI 87.31
87.12
87.59
Figure 4: Heatmaps generated under different model settings. (a) illustrates the attention map computed by vanilla Self-Attention. (b) illustrates ECCE without encoding local emotion context. (c) illustrates ECCE generated through complete two-stage encoding.
Table 2: Ablation results for Inverted Teacher-Student framework. We report WAF scores(%), and higher WAF indicates better performance. The Decline metric is obtained by calculating the difference in performance between 0% and 70% missing rates.
Dataset Training Framework Teacher Student Missing Rate
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 Decline
IEMOCAP (four-class) GCNet TF - 73.89 72.86 71.52 70.53 69.15 68.65 67.62 65.79 8.10
ITS TF TF 77.28 76.26 75.59 74.76 74.00 73.39 72.54 72.06 5.22
ITS TF NAS 77.47 76.87 76.44 76.02 75.48 74.72 74.19 73.84 3.63
TS NAS TF 76.15 75.48 74.70 74.00 73.37 72.86 72.33 71.77 4.38
IEMOCAP (six-class) GCNet TF - 56.98 56.02 54.79 53.11 51.67 49.84 47.85 47.46 9.52
ITS TF TF 58.59 58.30 57.29 56.79 55.89 55.18 54.74 54.31 4.28
ITS TF NAS 59.02 58.69 58.53 57.95 57.55 56.75 56.03 55.94 3.08
TS NAS TF 58.13 57.11 56.41 55.73 55.27 54.50 53.90 53.16 4.97
CMUMOSI GCNet TF - 84.99 81.65 75.95 75.99 72.27 69.11 62.79 64.66 20.33
ITS TF TF 85.93 82.45 79.49 77.21 72.97 69.46 65.68 66.88 19.05
ITS TF NAS 84.90 83.35 81.04 77.90 76.99 71.94 70.27 68.93 15.97
TS NAS TF 85.53 81.04 78.31 75.45 72.24 66.26 62.58 62.18 23.35
CMUMOSEI GCNet TF - 86.81 85.64 85.01 83.80 82.93 81.89 80.15 79.50 7.31
ITS TF TF 87.16 85.86 84.80 83.87 83.12 81.53 80.63 79.42 7.74
ITS TF NAS 86.71 86.40 84.97 84.40 83.25 82.00 80.65 80.21 6.50
TS NAS TF 86.72 86.22 84.35 84.06 82.52 82.02 80.10 79.04 7.68
Table 3: Comparison of classification performance under Independent Model Evaluation (IME). We present WAF scores (%).
Dataset Method Missing Rate
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 Average
IEMOCAP (four-class) CPM-Net 58.00 55.29 53.65 52.52 51.01 49.09 47.38 44.76 51.46
AE 74.82 71.36 67.40 62.02 57.24 50.56 43.04 39.86 58.29
CRA 76.26 71.28 67.34 62.24 57.04 49.86 43.22 38.56 58.23
MMIN 74.94 71.84 69.36 66.34 63.30 60.54 57.52 55.44 64.91
GCNet 78.36 77.48 77.34 76.22 75.14 73.80 71.88 71.38 75.20
ITEACNet 77.47 76.87 76.44 76.02 75.48 74.72 74.19 73.84 75.82
IEMOCAP (six-class) CPM-Net 41.05 37.33 36.22 35.73 35.11 33.64 32.26 31.25 35.32
AE 56.76 52.82 48.66 42.26 35.18 29.12 25.08 23.18 39.13
CRA 58.68 53.50 49.76 45.88 39.94 32.88 28.08 26.16 41.86
MMIN 56.96 53.94 51.46 48.42 45.60 42.82 40.18 37.84 47.15
GCNet 58.64 58.50 57.64 57.08 56.12 54.40 53.60 53.46 56.18
ITEACNet 59.02 58.69 58.53 57.95 57.55 56.75 56.03 55.94 57.56
CMUMOSI CPM-Net 71.90 68.91 71.12 70.59 67.95 65.88 64.02 61.79 67.77
AE 85.86 82.27 78.43 74.00 69.83 66.62 60.22 55.64 71.61
CRA 85.68 82.61 78.53 75.12 70.20 67.40 62.40 59.40 72.67
MMIN 85.20 81.91 78.22 74.60 70.14 67.72 64.04 61.53 72.92
GCNet 85.01 82.54 80.17 78.54 76.48 73.45 69.46 68.35 76.75
ITEACNet 84.90 83.35 81.04 77.90 76.99 71.94 70.27 68.93 76.92
CMUMOSEI CPM-Net 78.47 74.79 74.48 73.81 72.39 70.43 68.73 67.07 72.52
AE 86.66 84.37 82.58 80.57 78.80 76.43 74.26 72.81 79.56
CRA 86.48 84.19 82.25 80.12 78.55 75.85 74.07 72.46 79.25
MMIN 85.78 83.77 81.85 79.77 77.63 75.36 72.95 71.18 78.54
GCNet 87.12 86.50 85.50 84.53 83.55 82.44 80.27 80.20 83.76
ITEACNet 86.71 86.40 84.97 84.40 83.25 82.00 80.65 80.21 83.57
Evaluation Method
When evaluating the robustness of a model, it is crucial to consider its performance across various missing scenarios. Previous studies would define different scenarios, assuming that incidental factors hindering data collection would impact data at a fixed rate (Peng et al. 2020; Lian et al. 2023; Sun et al. 2023). Consequently, these works masked training and testing data at the same rate to simulate incomplete data scenarios, allowing them train the model and test its performance under this missing rate. This process will repeat with different missing rates to ensure experimental coverage of diverse missing scenarios. In this paper, we refer to this evaluation method as Independent Model Evaluation (IME).
However, in real-world scenarios, the aforementioned assumption is invalid, as the occurrence of incidental factors is uncertain. Therefore, the testing data that the model needs to handle will not maintain a fixed missing rate. To fill this gap, we propose a more realistic evaluation method called Unified Model Evaluation (UME). Once we obtain a well-trained model, we assume it will face a dynamic missing data scenario, where the missing rate of the testing data is no longer fixed.
Specifically, we mask the testing data at a fixed missing rate and then test the model’s performance under this missing rate without changing its weights. Subsequently, this process will repeat with different missing rates, and the average is taken as a comprehensive performance metric for the robustness.
Evaluation Metrics
Following (Lian et al. 2023), for CMU-MOSI and CMU-MOSEI, we use MAE as the loss function and select the model based on the best weighted average F1-score (WAF) on the validation set, reporting its WAF on the test set. For IEMOCAP, we adopt a 5-fold leave-one-session strategy and report the mean of the best WAF on the test sets. WAF is computed by excluding the samples with label 0.
Datasets
In this section, we provide a comprehensive overview of the dataset’s content and the data preprocessing methods employed in our study.
CMU-MOSI (Zadeh et al. 2016) contains 2,199 utterance-level video clips from 93 videos. Each video clip is annotated with an emotion intensity score ranging from -3 (strongly negative) to 3 (strongly positive).
CMU-MOSEI (Zadeh et al. 2018) is an iterative version of CMU-MOSI and includes 23,453 utterance-level video clips with the same annotation format as CMU-MOSI.
IEMOCAP (Busso et al. 2008) has two annotation formats: a four-class format and a six-class format. The four-class format consists of 5,531 utterance-level video clips, with each clip labeled as neutral, happy, sad, or angry. The six-class format includes 7,433 utterance-level video clips, with additional labels for excited and frustrated emotions.
Following (Lian et al. 2023)  [^1] https://github.com/zeroQiaoba/GCNet, we extract audio features using wav2vec (Schneider et al. 2019), text features using DeBERTa (Liu et al. 2019), and visual features using MA-Net (Zhao, Liu, and Wang 2021). Similar to (Lian et al. 2023), taking the case of 70% data missing as an example, when reading each mini-batch, each modality randomly resets 70% of the data to zero, but at least one modality is retained for each utterance.
Setup
We conduct experiments on all datasets using a unified setting, incorporating three modalities: text, audio, and video. The kernelsize of the $F_{local}$ is set to 7, and $dist_{m}$ is set to 30. Each Encoder Layer has a hidden dimension of 128, an FFN hidden layer ratio of 4.0, and Self-Attention with 8 head nums. The weights of Teacher model and Student model are optimized using AdaW with a learning rate of 5e-4, and Router are optimized using AdaW with a learning rate of 5e-3. The model is trained using a batch size of 32 for 100 epochs. We run each model three times and report average results. It’s worth noting that, in this paper, unless otherwise specified, the Teacher model is uniformly set to three Encoder Layers, while the Student model is uniformly set to four.
Results
Emotion Context Change Encoding Performance
Recognizing the significance of context changes in Conversational Emotion Recognition (CER), we propose a novel encoding method called Emotion Context Change Encoding (ECCE). This method takes into account context change information from both local and global perspectives. To comprehensively analyze the performance of ECCE, we compare the performance of teacher models under different settings, as ECCE is not influenced by data missing. The results are presented in Table 1, where ”Local” denotes the encoding of local emotion context, and ”Global” signifies the encoding of global context changes. As the local emotion context information cannot be directly utilized, we are unable to conduct independent performance testing on it.
Firstly, the incorporation of ECCE, derived from two-stage encoding involving Local and Global encoding, exhibits a consistent improvement in model performance across the four tasks compared to vanilla Self-Attention. Simultaneously, forgoing Local stage encoding and directly utilizing utterance features for encoding global context changes results in a decline in model performance across the four tasks.
Furthermore, as illustrated in Figure 4, we employ heatmaps for a more intuitive analysis of ECCE. Figure 4 depicts the attention map generated by vanilla Self-Attention, Figure 4 represents ECCE without Local stage encoding, and Figure 4 illustrates the complete ECCE. Darker colors indicate higher weights. All maps originate from the same conversation data, where both speakers initially expressed happiness. As the conversation progressed, a alternation between happy and excited emotions emerged.
As shown in Figure 4, the vanilla Self-Attention, when computing attention maps, tends to allocate higher weights to certain utterances that are more crucial throughout the entire conversation. However, it fails to consider the context in which the two utterances are situated and the changing between them. In contrast, the content in Figure 4 demonstrates a certain degree of local attention capability. However, this attention behavior appears somewhat ambiguous and does not accurately discern the emotional context.
Figure 4 illustrates the complete ECCE, which takes into account these aspects effectively. It can be observed that there is a strong local correlation in the top-left and bottom-right corners, indicating that the model can perceive significant differences in the changing context between the two parts of the conversation data. Additionally, there are three more pronounced local correlation regions in the bottom-right, corresponding to three distinct local contextual patterns: frequent emotional changes, a local context of happiness, and a local context of excitement. These results effectively demonstrate the advantage of ECCE in considering changes in local context and context change.
Table 4: Comparison of classification performance under Unified Model Evaluation (UME). We present WAF scores (%).
Dataset Method Missing Rate
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 Average
IEMOCAP (four-class) CPM-Net 44.05 44.33 42.63 41.60 41.18 41.48 41.21 41.37 42.23
AE 41.13 42.15 42.50 43.14 43.33 43.61 42.60 42.73 42.65
CRA 33.56 32.66 31.97 31.46 30.24 28.76 27.95 27.04 30.46
MMIN 70.80 69.25 67.25 64.93 62.32 59.60 56.93 54.21 63.16
GCNet 64.90 70.13 72.60 73.69 74.83 74.99 74.75 75.05 72.62
ITEACNet 77.13 76.46 76.08 75.67 74.50 73.88 73.17 72.78 74.96
IEMOCAP (four-class) CPM-Net 28.49 28.18 26.92 26.14 25.99 26.24 26.05 26.63 26.83
AE 24.60 24.41 24.55 24.04 23.87 23.05 23.23 22.63 23.80
CRA 16.16 23.69 23.14 22.63 21.94 20.98 19.73 19.36 20.95
MMIN 49.30 47.76 46.61 44.74 42.96 41.43 39.58 38.07 43.81
GCNet 44.39 49.31 52.88 54.24 55.20 54.42 54.52 54.21 52.40
ITEACNet 58.63 58.19 58.08 56.97 56.95 55.86 55.01 54.30 56.75
CMUMOSI CPM-Net 66.51 68.49 66.97 65.78 66.18 65.79 65.28 64.43 66.18
AE 62.36 60.37 59.58 57.57 56.49 53.62 51.56 50.03 56.45
CRA 61.27 58.54 55.78 54.24 50.90 48.16 45.92 41.80 52.08
MMIN 83.71 81.01 78.66 76.27 72.00 68.20 65.19 62.11 73.39
GCNet 68.79 68.17 70.47 70.63 68.59 69.47 67.88 69.55 69.19
ITEACNet 85.38 83.67 81.98 78.74 77.16 71.61 70.63 66.99 77.02
CMUMOSEI CPM-Net 74.17 74.58 74.51 74.06 72.83 71.32 69.67 68.26 72.43
AE 78.16 77.20 76.41 75.89 74.80 73.80 72.64 71.43 75.04
CRA 80.65 78.50 75.80 73.73 71.46 68.88 66.19 63.32 72.32
MMIN 85.24 83.02 80.30 78.08 74.90 72.08 67.86 65.85 75.91
GCNet 78.69 80.48 80.99 80.95 80.85 80.93 80.32 79.79 80.38
ITEACNet 87.08 86.15 85.38 84.43 83.17 81.79 80.92 79.90 83.60
Inverted Teacher-Student Traning Performance
Dealing with incomplete data poses a significantly greater challenge than handling complete data. However, current methods do not equip models for handling incomplete data with more stronger abilities to follow the performance of models trained with complete data. Therefore, we propose the Inverted Teacher-Student traning framework, which employs a student model with complex structures to handle incomplete data while emulating the performance of a teacher model with simpler structures trained with complete data. To comprehensively analyze the performance of ITS, we compare different training frameworks under different settings. The results are presented in Table 2, where the ”GCNet” training framework involves the teacher model receiving complete and incomplete data during the training phase, with gradients accumulated from both types of data and updated together at each step. The ”TS” training framework represents the vanilla Teacher-Student framework, where the teacher model is more complex than the student model. ”TF” is an abbreviation for Transformer. It’s worth noting that, in the TS training framework, the teacher model is equipped with four encoder layers, while the student model is equipped with three.
Firstly, the performance of the Transformer under GCNet training framework is not satisfactory. It not only fails to achieve outstanding performance with complete data but also exhibits poor capability in handling incomplete data. The IST training framework brings about a noticeable improvement in performance. The student model, guided by a simpler teacher model, not only performs better with complete data, surpassing the baselines across all four tasks, but also demonstrates significant performance improvements in handling incomplete data. In transitioning from complete data to a severe missing scenario with a 70% data loss, the ITS training framework exhibits less performance degradation. Compared to the GCNet training framework, ITS incurs a smaller loss of 2.88% in the four-class IEMOCAP task, 5.24% in the six-class IEMOCAP task, and 1.28% in the CMUMOSI task. While it shows a slight disadvantage in the CMUMOSEI task, the difference is negligible due to the similar performance under various missing rates for both frameworks.
Furthermore, we propose the efficient augmentation of the student model’s complexity using Neural Architecture Search (NAS), aiming to enhance its capability in handling incomplete data. While the introduction of NAS doesn’t necessarily result in superior performance on complete data, it significantly enhances the model’s capability to handle incomplete data. The student model enhanced through NAS achieved the smallest performance decline across all four tasks, with reductions of 3.63% in the four-class IEMOCAP, 3.08% in the six-class IEMOCAP, 15.97% in CMUMOSI, and 6.50% in CMUMOSEI. This substantial surpasses the previous settings. These results effectively demonstrate that by strengthening the computational capacity of the student model in ITS, it can handle incomplete data more effectively, even when trained with the same teacher.
Finally, we present additional results showcasing the performance of a simple student model under the guidance of a complex teacher model. The results indicate that a simple student model struggles to effectively acquire the knowledge from the teacher model, resulting in a more pronounced decline in performance. This comparison highlights the disparity between the ITS and vanilla teacher-student training frameworks in incomplete learning tasks.
Table 5: Model Params
Model Module Parameters
CPM-Net - 0.57M
AE - 1.29M
CRA - 2.11M
MMIN - 3.87M
GCNet - 34.14M
ITEACNet Teacher 2.19M
Student 3.88M
ITEACNet Performance
Independent Model Evaluation Performance
To demonstrate the outstanding performance of ITEACNet in CER tasks, we maintained the same evaluation method as previous works in Table 3, enabling a fair comparison with the baselines.
Firstly, the Average metric reflect the overall performance of the model under different missing scenarios. ITEACNet surpasses the current state-of-the-art (sota) work, GCNet, by 0.52%, 1.38%, and 0.17% in the four-class IEMOCAP, six-class IEMOCAP, and CMUMOSI tasks, respectively, achieving the new sota performance. Although ITEACNet’s performance in the CMUMOSEI task is 0.19% lower than GCNet, considering the carefully design of GCNet’s network structure, this performance remains highly competitive.
Furthermore, observing ITEACNet’s performance under different missing rates, we find that despite GCNet utilizing complex network structures to additionally process speaker information in conversation, there is still no significant difference between our method and GCNet across all four tasks. Even in severe missing scenarios, such as at a 70% missing rate, our method demonstrates more pronounced advantages in the four-class IEMOCAP and six-class IEMOCAP tasks, achieving performance improvements of 2.46% and 2.48%, respectively. These results collectively demonstrate the excellent robustness of ITEACNet.
Unified Model Evaluation Performance
To more accurately demonstrate the model’s robustness in real-world scenarios, we further employ the UME method to evaluate their performance.. The experimental results are presented in Table 4. As previous works did not provide experimental results for this evaluation method, all results presented here are reproduced by us. The parameter quantities for each model during the reproduction process are listed in Table 5.
Firstly, under the Average metric, ITEACNet exhibits a significant performance advantage across all four tasks. Compared to the state-of-the-art methods, our approach achieves performance improvements of 2.34%, 4.35%, 7.83%, and 3.22% in the four-class IEMOCAP, six-class IEMOCAP, CMUMOSI, and CMUMOSEI tasks, respectively.
Furthermore, observing ITEACNet’s performance under different missing rates reveals that despite the student model being exposed only to incomplete data during training, it still maintains excellent handling capabilities for complete data. This is attributed to the knowledge transferred from the teacher model, which processes complete data. In contrast, most baseline methods, except for the distillation-based MMIN, lack the ability to handle complete data, resulting in a substantial decrease in Average metric. Simultaneously, the student model in ITEACNet, utilizing an equivalent parameter quantity, achieves significantly superior performance compared to MMIN. This further underscores the superiority of our method.
Conclusion
In this paper, we propose a novel framework, ITEACNet, for incomplete multimodal learning in CER. ITEACNet has a new ECCE module to capture emotional context changes, and experimental results verify the significance of context changes in CER. It also includes a novel ITS training framework, allowing a complex student model to leverage incomplete data, following the performance of a simple teacher model trained on complete data. This design provides the student model with sufficient capability to handle incomplete data. We employ NAS to enhance the complexity of the student model, demonstrating the importance of a complex student model in incomplete learning. Finally, using an evaluation method closer to real-world scenarios, we show that previous methods lose the ability to handle complete data when dealing with incomplete data, while highlighting the importance of the ITS framework in incomplete learning.